OpenStudy (anonymous):

Solve each equation for x: 13x+4=5x+12

OpenStudy (shiraz14):

13x - 5x = 12 - 4 8x = 8 x = 1

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (shiraz14):

You're welcome